Read the two articles, entitled “The Clippers are Worth Nowhere Near What Ballmer is Paying” and “Why The Clippers are Worth $2 Billion”, and respond to the following questions. Where there is incomplete information, provide reasonable assumptions. 1. We are told the Clippers revenues at the time Steve Ballmer bought them was $128 million, and its profit was $15 million. a) At a purchase price of $2 billion, what is Ballmer’s ROA? b) To what level would the Clippers revenues need to rise in order for Ballmer to generate a ROA of 5%? 10%?

2. Assume the Clippers fixed costs (including things like Player salaries, administration and facilities costs) are equal to $100 million. Further assume that the Clippers TV and sponsorship revenue is $75 million and its only other source of revenue is ticket sales. a) How many tickets would the Clippers need to sell in order to breakeven, if the average contribution margin per ticket is $25? b) How many tickets would the Clippers need to sell in order to breakeven, if the average contribution margin per ticket is $50? c) The Clippers play 41 home games a season in a 20,000 seat arena. What is the likelihood of breaking even under scenarios a) and b) above? d) Are there any other variables that have not been contemplated above that could impact your response to c) above?
Answer #1

Q - 1

Part (a)

ROA = Profits / Total asset = Profits / total consideration = 15 / 2,000 = 0.75%

Part (b)

Profit margin = Profit / revenue = 15 / 128 = 11.72%

For ROA of 5%, desired revenue = ROA x total consideration / Profit margin = 5% x 2,000 / 11.72% = $ 853.33
million

For ROA of 10%, desired revenue = ROA x total consideration / Profit margin = 10% x 2,000 / 11.72% = $ 1,706.67
million

Q - 2

a) Tickets Clippers need to sell in order to breakeven = (Fixed costs - TV and sponsorship revenue) / Average contribution margin per ticket = (100 million - 75 million) / $ 25 = 1 million = 1,000,000

b) Tickets would the Clippers need to sell in order to breakeven= (Fixed costs - TV and sponsorship revenue) / Average contribution margin per ticket = (100 million - 75 million) / $ 50 = 0.5 million = 500,000

c) The Clippers play 41 home games a season in a 20,000 seat arena. What is the likelihood of breaking even under scenarios a) and b) above?

41 games x 20,000 seats = 820,000 seats < 1,000,000

Even if we assume 100% occupancy, Clippers will not break even under scenario (a)

820,000 seats > 500,000

Break even occupancy rate = 500,000 / 820,000 = 60.98% (a reasonable figure)

There appears reasonable likelihood of breaking even under scenarios (b) above

d) Are there any other variables that have not been contemplated above that could impact your response to c) above?

Yes, occupancy rate of the seats have not been considered. It's an important factor and could impact the response to (c) above.